OpenStudy (anonymous):

Solve x2 + 15x = -56. x = 8 and x = 7 x = -8 and x = -7 x = -8 and x = 7 x = 8 and x = -7

pooja195 (pooja195):

We need to factor this we need it to look like this \[\huge~\rm~ax^2+bx+c=0\]

pooja195 (pooja195):

how would we make it look like that ^

OpenStudy (solomonzelman):

test the discriminant \(\large\color{black}{ \displaystyle x^2 + 15x +56=0}\) \(\large\color{black}{ \displaystyle {\rm Discriminant} = \sqrt{15^2-4\cdot1\cdot56}}\) \(\large\color{black}{ \displaystyle {\rm Discriminant} = \sqrt{225-224}=1}\)

OpenStudy (solomonzelman):

u can factor it. Or go by the quadratic formula if you want to continue and not to lose my work